OpenStudy (anonymous):

Which problem situation below is represented by the equation below? y = 6x + 30 a. If a dog sitter charges $6 per hour, how much would she earn for x hours? b. If it takes six hours to earn $30, how long will it take to earn x dollars? c. If a landscape company charges $30 plus $6 per hour to mow a yard, how much would they charge for a job that takes x hours? d. If it costs $6 to travel one mile, how much will it cost to travel x miles?

jimthompson5910 (jim_thompson5910):

which one do you think it is?

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

b, I think it's b, but I really need a good grade. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

notice how b doesn't mention anything about a 6

jimthompson5910 (jim_thompson5910):

oh nvm i see the 6 hours

jimthompson5910 (jim_thompson5910):

but it's not B though

OpenStudy (anonymous):

oh, okay. :( my bad.

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

don't worry about being right or wrong

OpenStudy (anonymous):

okay. then which one is it? :/

jimthompson5910 (jim_thompson5910):

notice how choice C is If a landscape company charges $30 plus $6 per hour to mow a yard, how much would they charge for a job that takes x hours? this basically says that the fixed cost is $30 and then you add on a variable cost of $6 an hour so 30 + 6*(# of hours) = 30 + 6x =6x + 30

OpenStudy (anonymous):

omg, I was writing "my second guess is c!!!" ^.^

jimthompson5910 (jim_thompson5910):

lol beat you to it, but do you see how it's C?

OpenStudy (anonymous):

uh-huh! plus 30 and stuff, I see!!! ^.^

jimthompson5910 (jim_thompson5910):

ok that's great

OpenStudy (anonymous):

yes it is, awesome! Thank you! v.v

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

:)